OpenStudy (dls):

Derivative help! If f(x)=cos(pi/2 [x]-x^3),1

OpenStudy (dls):

\[\Huge f(x)=\cos(\frac{\pi}{2}[x]-x^3),1<x<2\]

OpenStudy (dls):

\[\LARGE and~[x]=GINT \le x ,then~f'(\sqrt[3]{\frac{\pi}{2}})=?\]

OpenStudy (dls):

@oldrin.bataku @Amber101 @mathslover @terenzreignz @Rayvenn.14

OpenStudy (dls):

@amistre64

OpenStudy (anonymous):

Is [ ] referring to round, ceil, or floor function? @DLS

OpenStudy (dls):

floor

OpenStudy (anonymous):

Notice that the floor function [x] for 1 < x < 2, will always be equal to 1 hence mimicking the horizontal line x = 1. So the derivative with the Chain Rule becomes:\[\bf f'(x)=3x^2\sin \left( \frac{\pi}{2}-x^3 \right)\]

OpenStudy (anonymous):

Notice that the [x] function will always equal to 1 for all 1 < x < 2. Hence we can eliminate it from the function itself and differentiate normally with the Chain Rule.

OpenStudy (dls):

yo!

OpenStudy (anonymous):

Now that we know the derivative, evaluate the derivative at the given x-value:\[\bf f'\left( \sqrt[3]{\frac{\pi}{2}} \right)=3\left( \sqrt[3]{\frac{\pi}{2}} \right)^2\sin\left( \frac{\pi}{2}-\frac{\pi}{2} \right)=0\] @DLS

OpenStudy (anonymous):

@DLS Do you understand?

OpenStudy (dls):

yes! thanks :D

OpenStudy (anonymous):

$$f(x)=\cos\left(\frac\pi2\lfloor x\rfloor-x^3\right)$$Given \(1<x<2\), note that \(\lfloor x\rfloor=1\) and therefore tangent lines for the floor will be flat: